### v2

@@ 4 Simulation Results
In this section, simulation experiments using the vehicle target state estimation are presented to illustrate the effectiveness of the proposed algorithm. Consider the following moving vehicle model(Zhou et al., 2022),
@@
This simulation considers 2 kinds of sensor network scenarios to validate the performance of the proposed algorithm:
  1. 1.
@@
The process noise $\omega$ and measurement noise $\nu_{i}$ are set with parameters $Q=0.5I$ and $R_{i}=0.5I$, respectively. Additionally, both the process noise and the measurement noise are bounded, i.e., $||\omega(k)||_{\infty}\leq 0.05$ and $||\nu_{i}(k)||_{\infty}\leq 0.05$ at all times. The observability matrices for each sensor are given by:
@@
In this simulation, only the velocity estimate $\hat{x}_{0,j}(k)$ in the sensor’s estimated results is tampered with by the attacker, while the position estimate remains unaffected. The attack signals are classified into two types: one type corresponds to a unstealthy attack, which exhibits large magnitudes most of the time; the other type corresponds to a stealthy attack, whose magnitude is close to the estimation noise most of the time. To better simulate different attack scenarios, the attack signals are designed based on the magnitude of the estimation noise, ensuring that stealthy and unstealthy attacks do not appear in the same simulation experiment. This setup meets the prerequisites of Lemma 2 shown in the APPENDIX A.
@@
Next, we compare the distribution of inter-subset gain mutual influence and intra-subset correlation under different partitioning strategies. Three partitioning strategies are considered: the partitioning strategy based on minimizing inter-subset gain mutual influence (Partitioning Strategy $1$), the suboptimal sensor partitioning strategy based on the Grassmann distance (Partitioning Strategy $2$), and the random division of $100$ sensors into $4$ groups (Partitioning Strategy $3$). For Partitioning Strategy $1$, the observable spaces of $C_{1}$ ($C_{2}$) and $C_{3}$ ($C_{4}$) overlap. Consequently, the partitioning result is that subset $1$ includes all sensors with measurement matrices $C_{1}$ and $C_{3}$, while subset $2$ includes all sensors with measurement matrices $C_{2}$ and $C_{4}$. As shown in TABLE 1, the inter-subset benefit influence is 00, while the intra-subset correlation is $0.5$, indicating that this strategy minimizes the inter-subset gain mutual influence. For Partitioning Strategy $2$, sensors are fully correlated only when their measurement matrices are identical. Thus, the partitioning result consists of 4 subsets, each containing one type of sensor. As shown in TABLE 1, the inter-subset gain mutual influence is $625$, while the intra-subset correlation is $1$, indicating that this strategy maximizes intra-subset correlation. For Partitioning Strategy $3$, by averaging the results of $100$ completely independent partitioning instances, we obtain the results shown in TABLE 1, where neither the inter-subset gain mutual influence nor the intra-subset correlation is optimal. Therefore, the proposed suboptimal partitioning strategy achieves the goal of minimizing inter-subset gain mutual influence while maximizing intra-subset correlation.
Table 1: Inter-subset Interaction and intra-subset Correlation under different Partitioning Strategy
@@
According to Theorem 1, the gain interactions only exist between $C_{2}$ and $C_{4}$, that is, under these two gain update strategies, the error in the distribution ratio vector only appears in subset $2$. The error curves are shown in Figure 1. To make the error curves smoother and more intuitive, the evaluation metric used is windowed RMSEs (W-RMSEs), with a window size of $10$ moments. It can be observed that the distribution ratio vector error in subset $2$ is bounded, with an average value of only $042.2871e-$ 04 over the entire time period, while the errors in subsets $1$, $3$, and $4$ remain zero throughout, which is consistent with the conclusion of Theorem 1.
@@
Scenario 2: Next, this simulation verifies that the proposed algorithm can ensure the secure state estimation of the system. Consider a complex distributed network scenario in which $500$ sensors (labeled $1-500$) are randomly deployed in a $200$m ×$\times 200$m area. Each sensor has the capability to communicate with sensors within a $30$m radius, forming an undirected graph. Similar to Scenario $1$, each sensor can independently measure the state of the vehicles, and the measurement matrix is classified into the $4$ types given in equation (4). First, measurement matrices are randomly assigned to the $500$ sensors, and the number of each type of sensor in the neighbor set of each sensor is recorded. The statistical results indicate that, for almost all sensors, their neighbor sets fail to satisfy the requirement stated in Remark 3, which mandates an equal number of sensors observing different state spaces. In this case, based on the set partitioning result, the communication direction between sensors is dynamically adjusted to control the distribution of neighbor sensors of each sensor. Ultimately, the entire network is transformed into a directed graph, ensuring that the neighbor set of each sensor satisfies Remark 3. Therefore, based on the proposed Algorithm 3.1 or Algorithm 3.2, the neighbor set of each sensor can be divided into $4$ subsets.
@@
At each moment $k=1:100$, the attacker randomly selects no more than half of the target sensors from the set of neighboring sensors of each sensor and launches stealthy attacks on their communication links. Therefore, the estimation results transmitted between sensors may be affected by FDIAs. Based on the D-ADS Algorithm, each sensor can preliminarily exclude a suspicious sensor set before fusing the information from neighboring sensors. This simulation compares the security of the system under different cases, including the case without attack detection, the case using the proposed D-ADS algorithm, and the case without any attack. For the no-attack scenario, the mean of the estimation errors of all sensors at each moment is used to calculate the RMSE. In contrast, for the case where attacks exist, considering that the security of the system can be intuitively described by the maximum estimation error, this simulation calculates the RMSE using the maximum estimation error at each moment. Figure 2 presents the RMSE curves under these cases, where there are three curves in total: the no-attack case (red), the proposed D-ADS algorithm detection (blue), and the case without attack detection (green). It can be seen that the blue curve is significantly lower than the green curve, indicating that the proposed algorithm ensures system security. However, there is still a certain gap between the blue and red curves, as stealthy attacks are challenging to detect accurately. This observation is consistent with the simulation results in Scenario $1$.
@@ 5 Conclusions
This paper investigates the performance of distributed attack detection scheduling algorithms in large-scale IoT networks. By analyzing the gain mutual influence between sensor subsets, an set partitioning strategy based on Grassmann distance is proposed, effectively reducing inter-subset mutual influence while enhancing intra-subset correlation. Theoretical analysis and simulation results demonstrate that the proposed strategy not only significantly narrows the performance gap between the D-ADS and ADS algorithm but also reduces computational cost. Future research can explore more advanced feature extraction methods to better capture complex information within the network, thereby further improving the efficiency and accuracy of attack detection.
6 APPENDIX A: The balance of the attack strategy
@@
First, let’s analyze the expected number of attacked sensors in each subset. Since the set of attacked sensors changes randomly at each moment, we estimate the mathematical expectation of the total number of attacked sensors over the entire period using the Monte Carlo method. The probability of a sensor being selected as an attack target is $q_{i}/|\mathcal{N}_{i}|$. Therefore, the expected number of attacked sensors in each subset is given by:
$\mathbb{E}[q_{i,g_{i}}]=|\mathcal{N}_{i,g_{i}}|\times\left(\frac{q_{i}}{|\mathcal{N}_{i}|}\right)\approx\frac{|\mathcal{N}_{i}|}{m_{i}}\times\frac{q_{i}}{|\mathcal{N}_{i}|}=\left\lfloor\frac{q_{i}}{m_{i}}\right\rfloor.$
@@
Therefore, this paper redefines the average malicious disturbance power from the attacker’s perspective, rather than from the specific communication link’s perspective. For the $l$-th attacker surrounding sensor $i$, where $l=1:q_{i}$, the average malicious disturbance power $\phi_{i,l}$ of the injected attack signal $z_{i,l}$ is defined as
$\phi_{i,l}=\lim_{T\to\infty}\frac{1}{T}\sum_{k=1}^{T}\left(z_{i,l}(k)\right)^{2}.$ (28)
@@
Based on the previous analysis, if the attacked sensors are completely random at each moment, the expected number of attacked sensors in each subset is $\left\lfloor q_{i}/{m_{i}}\right\rfloor$. Based on Assumption 1, the ratio of the average malicious disturbance power of each subset to the total malicious disturbance power of all attacks is $\left\lfloor q_{i}/m_{i}\right\rfloor/q_{i}=1/m_{i}$. This indicates that the attack intensity in each subset is balanced. This completes the proof.
@@
Then, a basic numerical simulation is utilized to verify that the attacker satisfies Lemma 2. Based on the set partitioning strategy in Algorithm 3.2, the 100 sensors are divided into four subsets: sensors 1–25 form subset 1, sensors 26–50 form subset 2, sensors 51–75 form subset 3, and sensors 76–100 form subset 4. Without considering specific attack types, at each time step, the attacker randomly selects 40 sensors for attack. The number of attacked sensors in each subset is shown in Figure 3. During the entire time period, the average number of sensors attacked in each subset was 9.83, 9.97, 10.04, and 10.16, respectively, which is close to the ideal average of 10.
@@
Building on this, when specific attack types are considered, the attack intensity distribution for each subset is shown in Figure 4. The vertical axis represents the ratio of the attack intensity in each subset to the total attack signal intensity, which is ideally 0.25. It can be seen from Figure 4 that, the intensity ratios fluctuate around 0.25. Specifically, in Figure 4(a), the mean attack intensity ratios for each subset are 0.2497, 0.2500, 0.2504, and 0.2499, while in Figure 4(b), the mean ratios are 0.2507, 0.2490, 0.2499, and 0.2504.
Figure 3: The number of attacked sensors in each subset
(a) Unstealthy attack
(b) Stealthy attack
Figure 4: Attack intensity ratio under unstealthy and stealthy attacks
7 APPENDIX B: Distributed Attack Detection Scheduling Algorithm

### v3

@@ 4 Simulation Results
In this section, simulation experiments are conducted on Automated Guided Vehicle (AGV) operation monitoring within an IoT scenario to illustrate the effectiveness of the proposed algorithm (Vlachos et al., 2024). In modern smart factories and warehouses, a large number of distributed sensors are deployed to continuously monitor the global operating states of AGVs, including their positions and velocities, in order to support real-time traffic management and scheduling. The AGV dynamics can be described by the following discrete-time motion model (Zhou et al., 2022),
@@
This simulation considers 2 kinds of sensor network scenarios to validate the performance of the proposed algorithm
  1. 1)
@@
And all simulations were conducted on a personal laptop equipped with an Intel i7-13700H CPU, 32 GB of RAM, running MATLAB R2022b.
@@
The process noise $\omega$ and measurement noise $\nu_{i}$ are set with parameters $Q=0.5I$ and $R_{i}=0.5I$, respectively. Additionally, both the process noise and the measurement noise are bounded, i.e., $\omega(k)||_{\infty}\leq 0.05$ and $\nu_{i}(k)||_{\infty}\leq 0.05$ at all times. The observability matrices for each sensor are given by:
@@
In this simulation, only the velocity estimate $\hat{x}_{0,\,j}(k)$ in the sensor’s estimated results is tampered with by the attacker, while the position estimate remains unaffected. The considered attack signals includes two types: one type corresponds to a unstealthy attack, which exhibits large magnitudes most of the time; the other type corresponds to a stealthy attack, whose magnitude is close to the estimation noise most of the time. To better simulate different attack scenarios, the attack signals are designed based on the magnitude of the estimation noise, ensuring that stealthy and unstealthy attacks do not appear in the same simulation experiment. This setup meets the prerequisites of Lemma 2 shown in the APPENDIX A.
@@
Next, we compare the distribution of inter-subset gain mutual influence and intra-subset correlation under different partitioning strategies. Three partitioning strategies are considered: the set partitioning strategy based on minimizing inter-subset gain mutual influence (Partitioning Strategy $1$), the set partitioning strategy based on the Grassmann distance (Partitioning Strategy $2$), and the random division of $100$ sensors into $4$ groups (Partitioning Strategy $3$). For Partitioning Strategy $1$, the observable spaces of $C_{1}$ ($C_{2}$) and $C_{3}$ ($C_{4}$) overlap. Consequently, the partitioning result is that subset $1$ includes all sensors with measurement matrices $C_{1}$ and $C_{3}$, while subset $2$ includes all sensors with measurement matrices $C_{2}$ and $C_{4}$. As shown in TABLE 2, the inter-subset benefit influence is 0, while the intra-subset correlation is $0.5$, indicating that this strategy minimizes the inter-subset gain mutual influence. For Partitioning Strategy $2$, sensors are fully correlated only when their measurement matrices are identical. Thus, the partitioning result consists of 4 subsets, each containing one type of sensor. As shown in TABLE 2, the inter-subset gain mutual influence is $625$, while the intra-subset correlation is $1$, indicating that this strategy maximizes intra-subset correlation. For Partitioning Strategy $3$, by averaging the results of $100$ completely independent partitioning instances, we obtain the results shown in TABLE 2, where neither the inter-subset gain mutual influence nor the intra-subset correlation is optimal. Therefore, the proposed set partitioning strategy achieves the goal of minimizing inter-subset gain mutual influence while maximizing intra-subset correlation.
Table 2: The inter-subset interaction and intra-subset correlation under different partitioning strategy
@@
According to Theorem 1, the gain interactions only exist between $C_{2}$ and $C_{4}$, that is, under these two gain update strategies, the error in the distribution ratio vector only appears in subset $2$. The error curves are shown in Figure 2. To make the error curves smoother and more intuitive, the evaluation metric used is windowed RMSEs (W-RMSEs), with a window size of $10$ moments. It can be observed that the distribution ratio vector error in subset $2$ is bounded, with an average value of only 2.2871e042.2871e-04 over the entire time period, while the errors in subsets $1$, 33, and $4$ remain zero throughout, which is consistent with the conclusion of Theorem 1.
@@
To more comprehensively evaluate the performance of the proposed algorithm, the absolute differences between D-ADS and the baseline ADS under different sensor selection methods is compared. The evaluation metric used is the average optimization rate defined in Definition 4, which is extended to the distributed case.
Definition 4
(Average Optimization Rate) (Suo et al., 2024c) For the entire time period $T$, the average optimization rate of the $i$-th sensor is defined as the ratio of the objective function (6) value of the candidate set $\mathcal{A}_{i,k}$ (or $\mathcal{A}_{i,\,k,\,g_{i}}$) selected by Algorithm 3.3, which is defined as $\frac{1}{T}\sum_{k=1}^{T}{(f_{k}(\mathcal{A}_{i,k})/f_{k}(\mathcal{A}_{i,k}^{*}))}$ (or $\frac{1}{T}\sum_{k=1}^{T}{(f_{k}(\mathcal{A}_{i,\,k,\,g_{i}})/f_{k}(\mathcal{A}_{i,\,k,\,g_{i}}^{*}))}$). And the average optimization rate for all the subset $g_{i}=1:m_{i}$ can be defined as $\frac{1}{T\cdot m_{i}}\sum_{k=1}^{T}\sum_{g_{i}=1}^{m_{i}}{(f_{k}(\mathcal{A}_{i,\,k,\,g_{i}})/f_{k}(\mathcal{A}_{i,\,k,\,g_{i}}^{*}))}$.
Based on the set partitioning strategy proposed in this paper for partitioning $100$ sensors and applying gain update method $1$ described in Section 3.4 to update element gains, we compare the average optimization rate for different sensor selection methods across each subset and the overall set. Theorem 2 states that the theoretical lower bound on the average optimization rate for any subset under the proposed D-ADS algorithm is 11/e1-1/e.
The two sensor selection methods considered are the probabilistic selection method and the ranking selection method. The probabilistic selection method employs a probabilistic approach in the first stage, followed by a uniform random selection algorithm in the second stage, whereas the ranking selection method adopts a ranking-based approach in the first stage, followed by a uniform random selection algorithm in the second stage. As shown in Table 3, the average optimization rate for individual subsets is unstable, typically fluctuating around the overall set’s average optimization rate. This is because it is difficult to ensure that the attack intensity is perfectly balanced across all subsets. Moreover, regardless of whether it is a subset or the overall set, the ranking selection method always achieves a higher average optimization rate than the probabilistic selection method. As shown in Table 4, the absolute difference in average optimization rate between ADS and D-ADS shows that the performance of the ADS algorithm is almost always better than that of the D-ADS algorithm in different situations. However, the difference is very limited, with the maximum absolute difference being $1.648\%$, which is tolerable.
Additionally, comparing the malicious information selection performance of the proposed algorithm under unstealthy and stealthy attacks, it is evident that the average optimization rate under unstealthy attacks is always higher than that under stealthy attacks.
Finally, to demonstrate the generality of the proposed algorithm, we further investigate the combined impact of different subset numbers and attacked sensor dimensions (i.e., the number of attacked sensors) under the case of 100 sensors, and analyze their effects on both the average number of attacked subsets and the fraction of attacked subsets. Specifically, the number of subsets is set to $[2,4,5,10,20,30,40,50]$, and the attacked sensor dimensions vary from 1 to 100 with a step size of 5.
The simulation results are shown in Figure 3. As illustrated in Figure 3(a), the number of attacked subsets increases with both the number of subsets and the attacked dimensions. In Figure 3(b), the fraction of attacked subsets also increases as the attacked dimensions grow. In the previous simulation, 100 sensors were divided into 4 subsets with 40 attacked dimensions, leading to all subsets being affected. However, the results in Figure 3 suggest that dividing the sensors into more subsets can significantly reduce the fraction of attacked subsets. This implies that, by ensuring independence or weak correlations among subsets, the number of subsets requiring gain updates can be reduced, thereby lowering the overall computational cost.
For example, when the sensors are divided into 20 subsets, each containing 5 sensors, and all subsets are mutually independent or weakly correlated, with 40 attacked dimensions, on average only about 13 subsets are affected. According to the proposed gain update Method 1, it is then sufficient to update only those 13 subsets (85 sensors in total), while the remaining subsets remain untouched. Compared with the previous case of 4 subsets, this reduces the computational cost by approximately $15\%$.
Table 3: Average optimization rate of each subset and the total set with different sensor selection methods
Algorithm Set Selection Method Average Optimization Rate
Probability Ranking Unstealthy Stealthy
D-ADS Subset 1 $\checkmark\checkmark$ 0.654 0.482
$\checkmark\checkmark$ 0.817 0.830
Subset 2 $\checkmark\checkmark$ 0.763 0.672
$\checkmark\checkmark$ 0.932 0.782
Subset 3 $\checkmark\checkmark$ 0.713 0.717
$\checkmark\checkmark$ 0.882 0.899
Subset 4 $\checkmark\checkmark$ 0.821 0.658
$\checkmark\checkmark$ 0.920 0.703
Average Subset $\checkmark\checkmark$ 0.738 0.632
$\checkmark\checkmark$ 0.888 0.804
ADS Total Set $\checkmark\checkmark$ 0.745 0.622
$\checkmark\checkmark$ 0.893 0.812
Table 4: The absolute difference in average optimization rate between ADS and D-ADS algorithms in different situations
Unstealthy Stealthy
[…]
(a) Number of attacked subsets
(b) Fraction of attacked subsets
Figure 3: The combined impact of different subset numbers and attacked sensor dimensions
@@
Scenario 2: Next, this simulation verifies that the proposed algorithm can ensure the secure state estimation of the system. Consider a complex distributed network scenario in which $500$ sensors (labeled 15001-500) are randomly deployed in a 200200m $\times 200$m area. Each sensor has the capability to communicate with sensors within a 3030m radius, forming an undirected graph. Similar to Scenario $1$, each sensor can independently measure the state of the vehicles, and the measurement matrix is classified into the $4$ types given in equation (4). First, measurement matrices are randomly assigned to the $500$ sensors, and the number of each type of sensor in the neighbor set of each sensor is recorded. The statistical results indicate that, for almost all sensors, their neighbor sets fail to satisfy the requirement stated in Remark 3, which mandates an equal number of sensors observing different state spaces. In this case, based on the set partitioning result, the communication direction between sensors is dynamically adjusted to control the distribution of neighbor sensors of each sensor. Ultimately, the entire network is transformed into a directed graph, ensuring that the neighbor set of each sensor satisfies Remark 3. Therefore, based on the proposed Algorithm 3.1 or Algorithm 3.2, the neighbor set of each sensor can be divided into $4$ subsets.
@@
At each moment $k=1:100$, the attacker randomly selects no more than half of the target sensors from the set of neighboring sensors of each sensor and launches stealthy attacks on their communication links. Therefore, the estimation results transmitted between sensors may be affected by FDIAs. Based on the D-ADS Algorithm, each sensor can preliminarily exclude a malicious information set before fusing the information from neighboring sensors. This simulation compares the security of the system under different cases, including the case without attack detection, the case using the proposed D-ADS algorithm, and the case without any attack. For the no-attack scenario, the mean of the estimation errors of all sensors at each moment is used to calculate the RMSE. In contrast, for the case where attacks exist, considering that the security of the system can be intuitively described by the maximum estimation error, this simulation calculates the RMSE using the maximum estimation error at each moment. Figure 4 presents the RMSE curves under these cases, where there are three curves in total: the no-attack case (red), the proposed D-ADS algorithm detection (blue), and the case without attack detection (green). It can be seen that the blue curve is significantly lower than the green curve, indicating that the proposed algorithm ensures system security. However, there is still a certain gap between the blue and red curves, as stealthy attacks are challenging to detect accurately. This observation is consistent with the simulation results in Scenario $1$.
@@ 5 Conclusions
Facing the impact of the expanding scale of IoT sensor networks on the efficiency of malicious information selection tasks, this paper investigates the application potential of distributed algorithms in large-scale IoT from the perspective of set partitioning strategy. Theoretical analysis and simulation results demonstrate that the proposed set partitioning strategy effectively narrows the performance gap between distributed and centralized methods, while ensuring that computational cost decreases as the number of subsets increases. The findings indicate that the proposed algorithm serves as an efficient cost-reduction strategy for detection algorithms, enabling them to be more feasible for resource-constrained edge or gateway devices in the Iot, while preserving detection reliability, thereby providing an efficient and feasible security protection path for practical IoT systems. Future research will explore more advanced feature extraction methods to capture the complex and dynamic information in IoT networks, thereby addressing sophisticated attacks such as Advanced Persistent Threats (APTs) more effectively.
Appendix A APPENDIX A: The balance of the attack strategy
@@
First, let’s analyze the expected number of attacked sensors in each subset. Since the set of attacked sensors changes randomly at each moment, we estimate the mathematical expectation of the total number of attacked sensors over the entire period using the Monte Carlo method. The probability of a sensor being selected as an attack target is $q_{i}/|\mathcal{N}_{i}|$. Therefore, the expected number of attacked sensors in each subset is given by $\mathbb{E}[q_{i,\,g_{i}}]=|\mathcal{N}_{i,\,g_{i}}|\times\left(\frac{q_{i}}{|\mathcal{N}_{i}|}\right)\approx\frac{|\mathcal{N}_{i}|}{m_{i}}\times\frac{q_{i}}{|\mathcal{N}_{i}|}=\left\lfloor\frac{q_{i}}{m_{i}}\right\rfloor$.
@@
Therefore, this paper redefines the average malicious disturbance power from the attacker’s perspective, rather than from the specific communication link’s perspective. For the $l$-th attacker surrounding sensor $i$, where $l=1:q_{i}$, the average malicious disturbance power $\phi_{i,\,l}$ of the injected attack signal $z_{i,\,l}$ is defined as $\phi_{i,\,l}=\lim_{T\to\infty}\frac{1}{T}\sum_{k=1}^{T}\left(z_{i,\,l}(k)\right)^{2}$. Based on the previous analysis, if the attacked sensors are completely random at each moment, the expected number of attacked sensors in each subset is $\left\lfloor q_{i}/{m_{i}}\right\rfloor$. Based on Assumption 1, the ratio of the average malicious disturbance power of each subset to the total malicious disturbance power of all attacks is $\left\lfloor q_{i}/m_{i}\right\rfloor/q_{i}=1/m_{i}$. This indicates that the attack intensity in each subset is balanced. This completes the proof.
@@
Then, a basic numerical simulation is conducted to verify that the attacker satisfies Lemma 2. According to the set partitioning strategy in Algorithm 3.2, the 100 sensors are divided into four subsets, where sensors 1–25 form subset 1, sensors 26–50 form subset 2, sensors 51–75 form subset 3, and sensors 76–100 form subset 4. Without considering specific attack types, the attacker randomly selects 40 sensors at each time step. The statistical results over the entire time horizon are summarized in Table 5, showing that the average number of sensors attacked in each subset is close to the ideal value of 10.
@@
When specific attack types are considered, the attack intensity distribution across subsets can also be evaluated. As shown in Table 5, the ratio of the attack intensity in each subset to the total attack signal intensity ideally equals 0.25. The numerical results demonstrate that the observed intensity ratios fluctuate slightly around this theoretical value, for both unstealthy and stealthy attacks.
Table 5: Attack statistics across subsets compared with the ideal values
Subset 1 Subset 2 Subset 3 Subset 4 Ideal
Average number of attacked sensors 9.83 9.97 10.04 10.16 10
Mean attack intensity ratio of unstealthy attacks 0.2497 0.2500 0.2504 0.2499 0.25
Mean attack intensity ratio of stealthy attacks 0.2507 0.2490 0.2499 0.2504 0.25
Appendix B APPENDIX B: Distributed Attack Detection Scheduling Algorithm